OpenStudy (anonymous):

Can someone help me slove this please!! My rectangular cabbage patch has a total area of 110 sq. ft. Find its perimeter in terms of the width x

OpenStudy (radar):

lenght times width = 110 sq ft. length = 11 and width =10 ft Perimeter is 11 + 11 + 10 + 10 or 42 feet. Now that is one answer out of many for that problem as you have not provided any conditions.

OpenStudy (anonymous):

Thank You!!!

OpenStudy (radar):

Usually a another piece of info is provided so that only a specific answer will work. Like the length is 1 ft greater than the width. Was there some more info for this problem>

OpenStudy (anonymous):

No thats all that was given to me

OpenStudy (radar):

O.K and that part about the length being 1 ft longer than the width makes the answer complete.

OpenStudy (anonymous):

My rectangular cabbage patch has a total area of 80 sq. ft. Find its perimeter in terms of the width x.

OpenStudy (radar):

l(length)=80/w Perimeter P = 2w + 2l P =160/w + 2w